# CrackSCF: Lightweight Cascaded Fusion Network for Robust and Efficient Structural Crack Segmentation

### v4

@@ 3.4 Staircase Cascaded Fusion Module
The design of our SCFM structure is shown in Fig. 2. As a segmentation head, to enhance useful semantic features in the MFE and LDE branches and ensure effective semantic information is not buried, we proposed the lightweight Pixel Attention Guided Fusion (PAF) module. The structure of PAF is shown in Fig. 2. We denote the pixels in the feature maps of the two branches as $\vec{v_{c}}$ and $\vec{v_{d}}$, respectively. The output can be expressed as:
$Output=\sigma\vec{v_{c}}+(1-\sigma)\vec{v_{d}}$ (12)
@@
To reduce computational resource usage, all convolution operations in this module utilize LRDS blocks. As seen in Fig. 3, using original convolution operations in the SCFM results in FLOPs and Params reaching 73.36G and 9.71M, respectively. Thus, lightweighting these operations is crucial. Replacing all convolution operations with LRDS blocks reduces FLOPs and Params by 92.87% and 97.22%, respectively, demonstrating the effectiveness of LRDS in network lightweighting.
@@
$\begin{split}&F_{i}^{O}=C_{i}\left(F_{i}^{MFE},F_{i-1}^{O}\right)\\ &+\mu\left(P_{i}\left(F_{i}^{MFE},F_{i-1}^{O}\right)\right)\odot P_{i}\left(F_{i}^{MFE},F_{i-1}^{O}\right)\end{split}$ (13)
@@ 3.5 Loss Function
The discrepancy between predictions and actual values is quantified by the loss function; a smaller value for this loss suggests superior model performance. Previously, Binary Cross-Entropy (BCE) was often used, but it may not effectively distinguish the minority crack pixels in the image. Combining BCE with Dice loss addresses this issue. BCE improves the probability distribution for each pixel, while Dice loss maximizes the overlap between predictions and true labels, enhancing robustness against imbalanced data [22] .
@@
4 Experiments
In this section, we analyze the differences between our TUT dataset and other public datasets, describe the experimental setup, and compare methods and metrics. We then analyze results and model complexity, perform ablation experiments.
@@ 4.1 Dataset Description
No. Datasets Year Num(train/val/test) Resolution Percentage of cracked pixels Scenarios
1 CFD [23] 2016 118(76/11/31) 480 width and 320 height 1.60% bitumen
2 GAPS509 [24] 2017 509(307/100/102) 540 width and 640 height 1.19% bitumen
3 DeepCrack [12] 2019 537(300/100/247) 544 width and 384 height 3.54% bitumen, cement
4 Crack500 [3] 2019 3368(1896/348/1124) 640 width and 360 height 6.01% bitumen, cement
5 CrackMap [25] 2023 120(72/12/36) 256 width and 256 height 4.10% bitumen
6 TUT (Ours) 2024 1408(987/139/282) 640 width and 640 height 3.16% bitumen, cement, bricks, runways, metal, blades, pipelines, tiles
@@
The previous works has used dataset scenarios that are relatively simple, such as cracks in cement or bitumen pavements only. These datasets have clear backgrounds and minimal interference, which cannot comprehensively evaluate the performance of models. To tackle this problem, we created the TUT dataset to enable better segmentation across different scenarios. Detailed information and comparisons with other datasets are list in Table 1.
@@
The TUT dataset comprises 1270 crack images, along with a small number of 138 images sourced from the Internet. TUT dataset contains 900 images of bitumen and cement with irregular crack morphology but flat backgrounds, 142 images of bricks and tiles with regular crack morphology, 228 images of plastic runways with extremely complex cracks, and a total of 138 images of metals, generator blades, and underground pipelines with complex lighting and background disturbances. To ensure reliability in model training and testing, all images were manually annotated by several researchers to generate binary labels, and the best-annotated images were selected as the final labeled images. Unlike other datasets with clear backgrounds, the TUT dataset images have complex backgrounds and cracks of intricate shapes. This evaluates the model’s capability to segment cracks in challenging conditions.
Figure 5: Illustration of the various scenarios of the TUT dataset with the original images and Ground Truth.
@@
As list in Table 1, our TUT dataset contains eight image scenarios, far more than the 1-2 scenarios of other public datasets, this allows for a more comprehensive evaluation of the model’s performance in detecting cracks under various scenarios. Fig. 5 shows crack images in different scenarios and their binary labeled images. The crack pixel ratio in the TUT dataset is 3.16%, which is moderate. This ensures the model learns useful information without being hindered by too few crack pixels and prevents the model from overly relying on the crack pixel ratio when there are too many crack pixels.
@@
As seen in Fig. 5, bitumen, cement, bricks, and plastic runways have extremely complex cracks due to their unique material properties. Particularly in plastic runways, both coarse and fine cracks are present, testing the model’s generalization to different types of cracks. In images of tiles, metal materials, wind turbine blades, and underground pipelines, the crack shapes are relatively simple, but the backgrounds are very complex with various external interferences such as low light, irrelevant areas, and surface characteristics. This can easily result in false detections, thereby testing the model’s ability to accurately extract crack areas from complex environments.
@@ 4.2 Implementation Details
Table 2: Comparison of results between CrackSCF and other methods on the public dataset and the TUT dataset. Best results are bolded and second best results are underlined
Methods CFD GAPS509 Crack500
ODS OIS P R F1 mIoU ODS OIS P R F1 mIoU ODS OIS P R F1 mIoU
RIND 0.6858 0.7021 0.6456 0.7534 0.6954 0.7592 0.5505 0.6023 0.4959 0.6494 0.5623 0.6909 0.6469 0.6483 0.6998 0.7245 0.7119 0.7381
UCTNet 0.6587 0.7006 0.6701 0.7205 0.6944 0.7479 0.5835 0.6188 0.5709 0.6752 0.6187 0.7107 0.7147 0.7265 0.6887 0.7843 0.7302 0.7674
SFIAN 0.7040 0.7159 0.6657 0.7512 0.7058 0.7684 0.6087 0.6320 0.6061 0.6451 0.6250 0.7198 0.6977 0.7348 0.6983 0.7742 0.7343 0.7604
CTCS 0.7122 0.7179 0.6896 0.7438 0.7157 0.7738 0.6029 0.6075 0.5693 0.6804 0.6199 0.7165 0.6941 0.7059 0.6940 0.7748 0.7322 0.7591
DTrCNet 0.7079 0.7123 0.6774 0.7486 0.7112 0.7710 0.6211 0.6276 0.6277 0.6592 0.6431 0.7272 0.7012 0.7241 0.6527 0.8280 0.7357 0.7627
Crackmer 0.6751 0.6754 0.6370 0.7237 0.6776 0.7503 0.4593 0.4642 0.6144 0.4663 0.5302 0.6581 0.6933 0.7097 0.6985 0.7572 0.7267 0.7591
SimCrack 0.7152 0.7204 0.6883 0.7557 0.7204 0.7756 0.6408 0.6471 0.6171 0.6997 0.6558 0.7368 0.7127 0.7308 0.7093 0.7984 0.7516 0.7715
Ours 0.7335 0.7376 0.6967 0.7848 0.7383 0.7864 0.6492 0.6552 0.6459 0.6812 0.6631 0.7411 0.7226 0.7353 0.7213 0.7922 0.7551 0.7764
CrackMap DeepCrack TUT (Ours)
RIND 0.6745 0.6943 0.6023 0.7586 0.6699 0.7425 0.8087 0.8267 0.7896 0.8920 0.8377 0.8391 0.7531 0.7891 0.7872 0.7665 0.7767 0.8051
UCTNet 0.6104 0.6195 0.5653 0.6748 0.6152 0.7079 0.8357 0.8579 0.8217 0.8857 0.8525 0.8564 0.7569 0.7670 0.8201 0.7667 0.7925 0.8110
SFIAN 0.7200 0.7465 0.6715 0.7668 0.7160 0.7748 0.8616 0.8928 0.8549 0.8692 0.8620 0.8776 0.7290 0.7513 0.7715 0.7367 0.7537 0.7896
CTCS 0.7289 0.7373 0.6911 0.7669 0.7270 0.7785 0.8819 0.8904 0.9011 0.8895 0.8952 0.8925 0.7940 0.7996 0.8202 0.8195 0.8199 0.8301
DTrCNet 0.7328 0.7413 0.6912 0.7681 0.7276 0.7812 0.8473 0.8512 0.8905 0.8251 0.8566 0.8661 0.7987 0.8073 0.7972 0.8441 0.8202 0.8331
Crackmer 0.7395 0.7437 0.7229 0.7467 0.7346 0.7860 0.8712 0.8785 0.8946 0.8783 0.8864 0.8844 0.7429 0.7640 0.7501 0.7656 0.7578 0.7966
SimCrack 0.7559 0.7625 0.7380 0.7672 0.7523 0.7963 0.8570 0.8722 0.8984 0.8549 0.8761 0.8744 0.7984 0.8090 0.8051 0.8371 0.8208 0.8334
Ours 0.7595 0.7651 0.7390 0.7708 0.7546 0.7987 0.8914 0.8963 0.9147 0.9013 0.9079 0.9001 0.8202 0.8266 0.8282 0.8484 0.8382 0.8473
2) Comparison Methods: To evaluate the model performance, we compared the CrackSCF with seven classical and latest SOTA methods on different datasets. These methods include RIND [26], UCTNet [8], SFIAN [11], CTCS [9], DTrCNet [7], Crackmer [10] and Simcrack [27].
3) Evaluation Metric: To quantitatively evaluate the performance of our proposed crack segmentation model, we used six main metrics [11]: Precision (P), Recall (R), F1 Score ($F1=\frac{2PR}{P+R}$), mean Intersection over Union (mIoU), Optimal Dataset Scale (ODS), and Optimal Image Scale (OIS).
@@ 4.3 Comparison with the SOTA methods
1) Results on the our TUT dataset: Fig. 6 shows the visual results of all methods on the TUT dataset. We manually highlighted easily missed crack pixel areas with red boxes and misdetected non-crack pixel areas with yellow boxes. As list in Table 2, our method achieved excellent results on the TUT dataset, leading in all metrics. Specifically, it outperformed the second-best methods on ODS, OIS, P, R, F1, and mIoU by 2.73%, 2.18%, 0.98%, 0.51%, 2.12%, and 1.67%, respectively. SimCrack [27] achieved the second-best results on ODS, OIS, F1, and mIoU, but as seen in Fig. 6, its segmentation performance on some key details has room for improvement compared to our method.
@@
2) Results on other public datasets: The CFD, CrackMap and GAPS509 datasets are characterized by slender and fine-grained crack structures, testing the model’s sensitivity to subtle and narrow crack pixels. As listed in Table 2, our proposed CrackSCF achieved the best performance on the CFD and CrackMap datasets. On both datasets, CrackSCF outperforms Simcrack by an average of 1.97%, 9.89%, 2.85%, 1.83%, and 1.00% on ODS, OIS, P, F1, and mIoU respectively. Although SimCrack [27] ranked second on most metrics, its actual segmentation performance on fine crack pixels and crack tips was unsatisfactory, as shown in Fig. 6. Other methods such as UCTNet [8] and DTrCNet [7] often failed to distinguish crack regions from background noise or missed crucial fine-grained crack details.
On the more challenging GAPS509 dataset, our method also achieved the highest performance on all metrics except recall. SimCrack obtained a higher recall value of 0.6997, about 2.72% higher than our method, but its precision was significantly lower at 0.6171, leading to a lower F1 score overall. Our method maintained a superior balance between precision and recall, with improvements of 1.31%, 1.25%, 2.90%, 1.11%, and 0.45% on ODS, OIS, P, F1, and mIoU, respectively. Visual results in Fig. 6 further confirm that our method effectively suppresses false detections in irrelevant background areas while preserving accurate segmentation of slender cracks.
Figure 6: Visualization results of CrackSCF and other methods on TUT and other public dataset. Key areas are highlighted with red boxes to indicate easily missed crack pixels and yellow boxes to indicate easily misdetected non-crack areas.
Both DeepCrack and Crack500 datasets are characterized by relatively wide crack regions, testing the model’s capability to capture extensive crack areas. As listed in Table 2, our method achieved the best results across all metrics on the DeepCrack dataset, outperforming the second-best methods by 1.08%, 0.39%, 1.51%, 1.04%, 1.42%, and 0.85% in terms of ODS, OIS, P, R, F1, and mIoU. Although CTCS [9] also performed well, it fell short in preserving crack detail segmentation, as illustrated in Fig. 6. This demonstrates the advantage of our method in accurately segmenting both wide crack regions and their finer structures. On the Crack500 dataset, our method achieved leading performance in five of the six evaluation metrics: ODS, OIS, P, F1, and mIoU, surpassing the second-best methods by 1.11%, 0.07%, 1.69%, 0.47%, and 0.64%. While DTrCNet [7] achieved a higher recall, our method outperformed it in precision by a notable 10.51%, leading to the highest F1 score and indicating a superior balance between detecting true crack pixels and suppressing false positives. Furthermore, as shown in Fig. 6, our method produces fewer erroneous detections in irrelevant background areas and achieves more accurate and refined segmentation of complex crack structures compared to competing approaches.
Table 3: Comparison of the complexity of CrackSCF with other methods
@@
Table 4: Comparison of CrackSCF using LRDS blocks with networks using other convolutions
Method ODS OIS P R F1 mIoU FLOPs ↓ Params ↓ FPS ↑
Original Conv 0.8145 0.8193 0.8294 0.8398 0.8346 0.8437 90.06G 17.86M 23
DS Conv 0.8117 0.8166 0.8239 0.8364 0.8301 0.8417 23.91G 8.38M 35
Low-Rank Conv 0.8154 0.8220 0.8220 0.8451 0.8334 0.8441 21.38G 4.63M 30
LRDS Conv 0.8202 0.8266 0.8282 0.8484 0.8382 0.8473 9.26G 4.79M 36
Table 5: Comparison of SCFM performance with other segmentation heads
SegHead ODS OIS P R F1 mIoU FLOPs ↓ Params ↓ FPS ↑
PID 0.7120 0.7198 0.7921 0.6742 0.7284 0.7777 36.69G 5.44M 21
SegFormer 0.7446 0.7634 0.7533 0.7786 0.7658 0.7983 29.72G 6.20M 23
UNet 0.7836 0.7951 0.7829 0.8248 0.8033 0.8222 40.59G 12.34M 20
Hamburger 0.7447 0.7609 0.7075 0.8328 0.7670 0.7969 168.23G 7.45M 14
SCFM 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473 9.26G 4.79M 36
Table 6: Comparison of the performance of using our lightweight deformable attention in LDE with other variants
Attention ODS OIS P R F1 mIoU FLOPs Params FPS↑
Common MHA 0.7919 0.8009 0.7999 0.8240 0.8118 0.8279 15.98G 5.80M 16
MHA + LR_Linear 0.7924 0.8003 0.7988 0.8284 0.8134 0.8286 8.87G 4.77M 14
Common MSDA 0.8058 0.8151 0.8052 0.8467 0.8257 0.8373 16.37G 5.82M 28
MSDA + LR_Linear 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473 9.26G 4.79M 36

### v5

@@ 2.4 Staircase Cascaded Fusion Module
Figure 6: Illustration of the PAF. It effectively reduces required computing resources by embedding LRDS.
The design of our SCFM structure is shown in Figure 2. As a segmentation head, to enhance useful semantic features in the MFE and LDE branches and ensure effective semantic information is not buried, this paper proposed the lightweight Pixel Attention Guided Fusion (PAF) module. The structure of PAF is shown in Figure 6. We denote the pixels in the feature maps of the two branches as $\vec{v_{c}}$ and $\vec{v_{d}}$, respectively. The output can be expressed as:
$\sigma=\text{Sigmoid}(f_{c}(\vec{v_{c}})\cdot f_{d}(\vec{v_{d}}))$ (12)
where $f_{c}$ and $f_{d}$ denote pointwise convolution and batch normalization operations, respectively. Thus, the output of PAF can be represented as:
$\text{Output}=\sigma\vec{v_{c}}+(1-\sigma)\vec{v_{d}}$ (13)
@@
To reduce computational resource usage, all convolution operations in this module utilize LRDS blocks. As seen in Figure 3, using original convolution operations in the SCFM results in FLOPs and Params reaching 73.36G and 9.71M, respectively. Thus, lightweighting these operations is crucial. Replacing all convolution operations with LRDS blocks reduces FLOPs and Params by 92.87% and 97.22%, respectively, demonstrating the effectiveness of LRDS in network lightweighting.
@@
$\begin{split}&F_{i}^{\text{O}}=C_{\text{i}}\left(F_{i}^{\text{MFE}},F_{i-1}^{\text{O}}\right)\\ &+\mu\left(P_{i}\left(F_{i}^{\text{MFE}},F_{i-1}^{\text{O}}\right)\right)\odot P_{i}\left(F_{i}^{\text{MFE}},F_{i-1}^{\text{O}}\right)\end{split}$ (14)
@@ 2.5 Loss Function
The discrepancy between predictions and actual values is quantified by the loss function; a smaller value for this loss suggests superior model performance. Previously, Binary Cross-Entropy (BCE) [16] was often used, but it may not effectively distinguish the minority crack pixels in the image. Combining BCE with Dice loss [15] addresses this issue. BCE improves the probability distribution for each pixel, while Dice loss maximizes the overlap between predictions and true labels, enhancing robustness against imbalanced data.
@@
In this section, the paper analyzes the differences between the TUT dataset and other public datasets, describes the experimental setup, and compares various methods and metrics. Subsequently, the results and model complexity are analyzed, and ablation experiments are performed.
3 Experimental Results and Discussion
@@ 3.1 Dataset Description
No. Datasets Num Resolution Crack Pixels Scenarios
1 CFD [28] 118 480 $\times$ 320 1.60% 1
2 GAPS509 [9] 509 540 $\times$ 640 1.19% 1
3 DeepCrack [17] 537 544 $\times$ 384 3.54% 2
4 Crack500 [40] 3368 640 $\times$ 360 6.01% 2
5 CrackMap [14] 120 256 $\times$ 256 4.10% 1
6 AEL [40] 58 768 $\times$ 512 0.67% 1
7 CrackTree206 [46] 206 800 $\times$ 600 0.32% 1
8 LSCD [41] 143 400 $\times$ 400 3.18% 2
9 CrackSC [12] 197 320 $\times$ 480 2.02% 2
10 TUT (Ours) 1408 640 $\times$ 640 3.16% 8
Figure 7: Illustration of the images of each scenarios in the TUT dataset.
@@
The previous works has used dataset scenarios that are relatively simple, such as cracks in cement or bitumen pavements only. These datasets have clear backgrounds and minimal interference, which cannot comprehensively evaluate the performance of models. To tackle this, this paper created the TUT dataset to enable better segmentation across different scenarios. Detailed information and comparisons with other datasets are list in Table 1.
@@
The TUT dataset comprises 1270 crack images captured with our mobile phones, along with a small number of 138 images sourced from the Internet. TUT dataset contains 900 images of bitumen and cement with irregular crack morphology but flat backgrounds, 142 images of bricks and tiles with regular crack morphology, 228 images of plastic runways with extremely complex cracks, and a total of 138 images of metals, generator blades, and underground pipelines with complex lighting and background disturbances. To ensure reliability in model training and testing, all images were manually annotated by several researchers to generate binary labels, and the best-annotated images were selected as the final labeled images. Unlike other datasets with clear backgrounds, the TUT dataset images have complex, noisy backgrounds and cracks of intricate shapes. This evaluates the model’s capability to segment cracks in challenging conditions.
@@
As shown in Table 1, our TUT dataset contains eight image scenarios, far more than the 1-2 scenarios of other public datasets, this allows for a more comprehensive evaluation of the model’s performance in detecting cracks under various scenarios. Figure 7 shows crack images in different scenarios and their binary labeled images. The crack pixel ratio in the TUT dataset is 3.16%, which is moderate. This ensures the model learns useful information without being hindered by too few crack pixels and prevents the model from overly relying on the crack pixel ratio when there are too many crack pixels.
@@
As seen in Figure 7, bitumen, cement, bricks, and plastic runways have extremely complex cracks due to their unique material properties. Particularly in plastic runways, both coarse and fine cracks are present, testing the model’s generalization to different types of cracks. In images of tiles, metal materials, wind turbine blades, and underground pipelines, the crack shapes are relatively simple, but the backgrounds are very complex with various external interferences such as low light, irrelevant areas, and surface characteristics. This can easily result in false detections, thereby testing the model’s ability to accurately extract crack areas from complex environments.
@@ 3.2 Implementation Details
2) Comparison Methods: In evaluating model performance, we compared the CrackSCF network with 12 classic segmentation methods and the latest SOTA methods on different datasets. These methods include HED [37], UNet++ [44], Deeplabv3+ [2], AttuNet [22], RCF [19], RIND [24], UCTNet [33], SFIAN [4], CT-crackseg [30], DTrCNet [35], Crackmer [34] and Simcrack [13].
Figure 8: Visualization results of CrackSCF and other methods on public datasets. Key areas are highlighted with red boxes to indicate easily missed crack pixels and yellow boxes to indicate easily misdetected non-crack areas.
3) Evaluation Metric: To quantitatively evaluate the performance of our proposed crack segmentation model, we used six main metrics: Precision (P), Recall (R), F1 Score ($F1=\frac{2PR}{P+R}$), Mean Intersection over Union (mIoU), Optimal Dataset Scale (ODS), and Optimal Image Scale (OIS). Mean Intersection over Union (mIoU) measures the average ratio of the intersection and union of the predicted results and the true results. The calculation formula for mIoU is:
$mIoU=\frac{1}{i+1}\sum_{a=0}^{i}\frac{p_{aa}}{\sum_{b=0}^{i}p_{ab}+\sum_{b=0}^{i}p_{ba}-p_{aa}}$ (16)
where $i$ is the number of classes, $a$ denotes the true value, $b$ denotes the predicted value, and $p_{ab}$ represents the number of pixels predicted as $a$ for $b$. Given that our crack segmentation task involves binary classification, we set $i=1$.
Optimal Dataset Scale (ODS) indicates the optimal segmentation performance metric obtained when a fixed threshold $t$ is selected over the entire dataset. It is defined as:
$ODS=\max_{t}\frac{2\cdot P_{t}\cdot R_{t}}{P_{t}+R_{t}}$ (17)
Optimal Image Scale (OIS) indicates the optimal segmentation performance metric obtained when an optimal threshold $t$ is selected for each image individually. It is defined as:
$OIS=\frac{1}{N}\sum_{a=1}^{N}\max_{t}\frac{2\cdot P_{t,a}\cdot R_{t,a}}{P_{t,a}+R_{t,a}}$ (18)
where $N$ represents the total number of images.
@@ 3.3 Comparison with the SOTA methods
Our proposed method, CrackSCF, was compared with 12 other methods across five public datasets and our own TUT dataset.
Table 2: Comparison results on the CFD dataset. Best results are bolded and second best results are underlined.
Methods ODS OIS P R F1 mIoU
HED 0.6276 0.6494 0.5467 0.7374 0.6279 0.7240
UNet++ 0.6586 0.6836 0.5995 0.7608 0.6706 0.7432
DeeplabV3+ 0.6910 0.7084 0.6264 0.7730 0.6920 0.7709
AttuNet 0.6710 0.6870 0.6240 0.7220 0.6694 0.7576
RCF 0.6434 0.6624 0.5792 0.7348 0.6478 0.7343
RIND 0.6858 0.7021 0.6456 0.7534 0.6954 0.7592
UCTNet 0.6587 0.7006 0.6701 0.7205 0.6944 0.7479
SFIAN 0.7040 0.7159 0.6657 0.7512 0.7058 0.7684
CT-crackseg 0.7122 0.7179 0.6896 0.7438 0.7157 0.7738
DTrCNet 0.7079 0.7123 0.6774 0.7486 0.7112 0.7710
Crackmer 0.6751 0.6754 0.6370 0.7237 0.6776 0.7503
SimCrack 0.7152 0.7204 0.6883 0.7557 0.7204 0.7756
Ours 0.7335 0.7376 0.6967 0.7848 0.7383 0.7864
1) Results on the CFD dataset: The crack areas in the CFD dataset are very slender, testing the model’s sensitivity to fine-grained crack pixels. As shown in Table 2, our proposed CrackSCF attained the best results across all metrics for the CFD dataset. Specifically, on ODS, OIS, P, R, F1, and mIoU, it outperformed the second-best methods by 0.48%, 0.34%, 0.14%, 0.35%, 0.31%, and 0.30%, respectively. Notably, SimCrack [13] achieved the second-best performance on all metrics except for P and R, only behind our method. However, as seen in Figure 8, its actual segmentation performance is not satisfactory, while our method achieves high-quality segmentation results on various fine-grained crack pixels. Other methods, such as UCTNet [33] and DTrCNet [35], tend to misdetect background noise or miss key detail areas of crack pixels.
Table 3: Comparison results on the GAPS509 dataset. Best results are bolded and second best results are underlined.
Methods ODS OIS P R F1 mIoU
HED 0.4014 0.4509 0.4300 0.4624 0.4456 0.6282
UNet++ 0.6255 0.6484 0.6160 0.6684 0.6411 0.7300
DeeplabV3+ 0.6160 0.6425 0.6050 0.6669 0.6345 0.7378
AttuNet 0.5955 0.6174 0.5924 0.6519 0.6208 0.7247
RCF 0.4532 0.4817 0.4420 0.5428 0.4872 0.6472
RIND 0.5505 0.6023 0.4959 0.6494 0.5623 0.6909
UCTNet 0.5835 0.6188 0.5709 0.6752 0.6187 0.7107
SFIAN 0.6087 0.6320 0.6061 0.6451 0.6250 0.7198
CT-crackseg 0.6029 0.6075 0.5693 0.6804 0.6199 0.7165
DTrCNet 0.6211 0.6276 0.6277 0.6592 0.6431 0.7272
Crackmer 0.4593 0.4642 0.6144 0.4663 0.5302 0.6581
SimCrack 0.6408 0.6471 0.6171 0.6997 0.6558 0.7368
Ours 0.6492 0.6552 0.6459 0.6812 0.6631 0.7411
2) Results on the GAPS509 dataset: Unlike the CFD dataset, the GAPS509 dataset contains not only finer cracks but also some irrelevant background interference, testing the model’s ability to extract crack areas under high interference conditions. As shown in Table 3, our method achieved the highest performance on all metrics except R. Although SimCrack [13] achieved 0.6997 on the R metric, leading our method by about 2.72%, its P metric was 0.6171, far lower than our method. The highest F1 score obtained by our method indicates a good balance between P and R. Specifically, our method outperformed the second-best methods on ODS, OIS, P, F1, and mIoU by 1.31%, 1.25%, 2.90%, 1.11%, and 0.45%, respectively. Additionally, as seen in Figure 8, our method almost perfectly avoids irrelevant background noise, while most other methods falsely detect irrelevant backgrounds and perform worse in segmenting crack pixel areas compared to our method.
Table 4: Comparison results on the DeepCrack dataset. Best results are bolded and second best results are underlined.
Methods ODS OIS P R F1 mIoU
HED 0.7552 0.8237 0.7231 0.7678 0.7448 0.8007
UNet++ 0.8419 0.8893 0.8474 0.8533 0.8503 0.8621
DeeplabV3+ 0.7727 0.8354 0.7363 0.8323 0.7813 0.8565
AttuNet 0.8153 0.8379 0.8066 0.8069 0.8068 0.8578
RCF 0.7666 0.8376 0.7423 0.7611 0.7516 0.8110
RIND 0.8087 0.8267 0.7896 0.8920 0.8377 0.8391
UCTNet 0.8357 0.8579 0.8217 0.8857 0.8525 0.8564
SFIAN 0.8616 0.8928 0.8549 0.8692 0.8620 0.8776
CT-crackseg 0.8819 0.8904 0.9011 0.8895 0.8952 0.8925
DTrCNet 0.8473 0.8512 0.8905 0.8251 0.8566 0.8661
Crackmer 0.8712 0.8785 0.8946 0.8783 0.8864 0.8844
SimCrack 0.8570 0.8722 0.8984 0.8549 0.8761 0.8744
Ours 0.8914 0.8963 0.9147 0.9013 0.9079 0.9001
3) Results on the DeepCrack dataset: The cracks in the DeepCrack dataset are often relatively wide, testing the model’s ability to extract a large range of crack pixels. As shown in Table 4, our method achieved the highest results on all metrics for the DeepCrack dataset. Specifically, our method outperformed the second-best methods on ODS, OIS, P, R, F1, and mIoU by 1.08%, 0.39%, 1.51%, 1.04%, 1.42%, and 0.85%, respectively. Notably, CT-crackseg also achieved good results on this dataset, but as seen in Figure 8, its segmentation performance on detailed cracks is not as good as our method, indicating that our method can effectively handle crack details.
4) Results on the Crack500 dataset: Similar to the DeepCrack, the Crack500 dataset also has relatively wide crack areas, but its structure is complex, posing higher requirements on model performance. As shown in Table 5, our method achieved the best results on five metrics for the Crack500 dataset, except for R. It outperformed the second-best methods on ODS, OIS, P, F1, and mIoU by 1.11%, 0.07%, 1.69%, 0.47%, and 0.64%, respectively. Although DTrCNet [35] leads on the R metric, our method surpasses it on the P metric by 10.51%. The highest F1 score achieved by our method indicates a good balance between P and R. Additionally, as seen in Figure 8, our method does not produce false detections in irrelevant background areas, whereas other methods either produce false detections or lack detailed segmentation in crack areas.
Table 5: Comparison results on the Crack500 dataset. Best results are bolded and second best results are underlined.
Methods ODS OIS P R F1 mIoU
HED 0.6191 0.6537 0.6150 0.7281 0.6668 0.7140
UNet++ 0.6965 0.7274 0.6989 0.7805 0.7374 0.7605
DeeplabV3+ 0.6943 0.7211 0.6970 0.7569 0.7257 0.7746
AttuNet 0.6586 0.6890 0.6490 0.7608 0.7005 0.7543
RCF 0.6054 0.6361 0.6037 0.7136 0.6540 0.7068
RIND 0.6469 0.6483 0.6998 0.7245 0.7119 0.7381
UCTNet 0.7147 0.7265 0.6887 0.7843 0.7302 0.7674
SFIAN 0.6977 0.7348 0.6983 0.7742 0.7343 0.7604
CT-crackseg 0.6941 0.7059 0.6940 0.7748 0.7322 0.7591
DTrCNet 0.7012 0.7241 0.6527 0.8280 0.7357 0.7627
Crackmer 0.6933 0.7097 0.6985 0.7572 0.7267 0.7591
SimCrack 0.7127 0.7308 0.7093 0.7984 0.7516 0.7715
Ours 0.7226 0.7353 0.7213 0.7922 0.7551 0.7764
5) Results on the CrackMap dataset: Similar to the CFD dataset, the cracks in the CrackMap dataset are also relatively slender. As shown in Table 6, our method leads all other methods on all metrics, outperforming the second-best methods on ODS, OIS, P, R, F1, and mIoU by 0.48%, 0.34%, 0.14%, 0.35%, 0.31%, and 0.30%, respectively. Notably, SimCrack [13] achieved the second-best results on this dataset, but as seen in Figure 8, SimCrack’s segmentation performance on the crack tips is not satisfactory, whereas our method effectively identifies these pixels.
Table 6: Comparison results on the CrackMap dataset. Best results are bolded and second best results are underlined.
Methods ODS OIS P R F1 mIoU
HED 0.6022 0.6240 0.5281 0.7063 0.6043 0.7026
UNet++ 0.7030 0.7222 0.6773 0.7209 0.6984 0.7631
DeeplabV3+ 0.6617 0.6793 0.5853 0.7532 0.6587 0.7460
AttuNet 0.6618 0.6773 0.6233 0.6970 0.6581 0.7454
RCF 0.5552 0.5683 0.4590 0.7164 0.5595 0.6758
RIND 0.6745 0.6943 0.6023 0.7586 0.6699 0.7425
UCTNet 0.6104 0.6195 0.5653 0.6748 0.6152 0.7079
SFIAN 0.7200 0.7465 0.6715 0.7668 0.7160 0.7748
CT-crackseg 0.7289 0.7373 0.6911 0.7669 0.7270 0.7785
DTrCNet 0.7328 0.7413 0.6912 0.7681 0.7276 0.7812
Crackmer 0.7395 0.7437 0.7229 0.7467 0.7346 0.7860
SimCrack 0.7559 0.7625 0.7380 0.7672 0.7523 0.7963
Ours 0.7595 0.7651 0.7390 0.7708 0.7546 0.7987
Table 7: Comparison results on the TUT dataset. Best results are bolded and second best results are underlined.
Methods ODS OIS P R F1 mIoU
HED 0.6960 0.7254 0.7080 0.7431 0.7251 0.7683
UNet++ 0.7702 0.7948 0.7725 0.7974 0.7848 0.8160
DeeplabV3+ 0.7504 0.7777 0.7155 0.8231 0.7655 0.8177
AttuNet 0.7539 0.7795 0.7558 0.7945 0.7747 0.8204
RCF 0.6821 0.7138 0.7106 0.6877 0.6990 0.7598
RIND 0.7531 0.7891 0.7872 0.7665 0.7767 0.8051
UCTNet 0.7569 0.7670 0.8201 0.7667 0.7925 0.8110
SFIAN 0.7290 0.7513 0.7715 0.7367 0.7537 0.7896
CT-crackseg 0.7940 0.7996 0.8202 0.8195 0.8199 0.8301
DTrCNet 0.7987 0.8073 0.7972 0.8441 0.8202 0.8331
Crackmer 0.7429 0.7640 0.7501 0.7656 0.7578 0.7966
SimCrack 0.7984 0.8090 0.8051 0.8371 0.8208 0.8334
Ours 0.8202 0.8266 0.8282 0.8484 0.8382 0.8473
Figure 9: Visualization results of CrackSCF and other methods on TUT dataset. Key areas are highlighted with red boxes to indicate easily missed crack pixels and yellow boxes to indicate easily misdetected non-crack areas.
@@
6) Results on the TUT dataset: Figure 9 presents the visual results of all methods on the TUT dataset. Easily missed crack pixel areas are highlighted with red boxes, and misdetected non-crack pixel areas are indicated with yellow boxes. As shown in Table 7, the proposed method achieved excellent results on the TUT dataset, leading in all metrics. Specifically, it outperformed the second-best methods on ODS, OIS, P, R, F1, and mIoU by 2.73%, 2.18%, 0.98%, 0.51%, 2.12%, and 1.67%, respectively. While SimCrack [13] achieved the second-best results on ODS, OIS, F1, and mIoU, as observed in Figure 9, its segmentation performance on certain key details exhibits room for improvement compared to the proposed method.
@@
Table 8: Comparison of the complexity of CrackSCF with other methods. Best results are bolded and second best results are underlined.